# custom.bib

### v2

@@ Abstract
While Rotary Position Embeddings (RoPE) for large language models have become widely adopted, their application for other modalities has been slower. Here, we introduce Lie group Relative position Encodings (LieRE) that goes beyond RoPE in supporting n-dimensional inputs. We evaluate the performance of LieRE on 2D and 3D image classification tasks and observe that LieRE leads to marked relative improvements in performance (up to 9.7% for 2D and up to 25.5% for 3D), training efficiency (3.5x reduction), data efficiency (30%) compared to the baselines of DeiT III, RoPE-Mixed and Vision-Llama. https://github.com/Stanford-AIMI/LieRE
@@ 1 Introduction
While the attention mechanism has achieved widespread use, especially as part of the transformer architecture, attention is invariant to the order of its inputs and requires another mechanism to capture positional information of input tokens [vaswani2017attention]. This has spurred a line of work in the subarea of positional encodings—methods of encoding positional information in attention mechanisms.
In particular, Rotary Position Encoding (RoPE) has emerged as a technique for encoding relative positional information of text tokens in transformer-based models [RoFormer2024Su]. RoPE’s ability to capture relative position information has made it a popular choice for open-source language foundation models such as LLaMA. In particular, RoPE implicitly captures relative positions. When the token in position $p_{1}$ attends to a token in position $p_{2}$, the effect of RoPE depends on $p_{1}-p_{2}$.
Despite the success of RoPE in sequence tasks [touvron2023llama, chowdhery2023palm], it is designed for one-dimensional sequence data. This has resulted in slow adoption for modalities with higher dimensional data, such or data that includes a temporal dimension like videos.
Our work aims to answer whether a single position encoding scheme can work well across both 2D and 3D modalities. If possible, this would enable the use of a simpler common model backbone across a variety of tasks.
[…]
@@
Figure 1: LieRE sketch, where $A$ is a learnable skew symmetric matrix and $R_{i}=\exp(A[x\;y\;z])\in\mathbb{R}^{d\times d}$ is the rotation matrix for the first patch in the top left corner of the input with the position $P_{i}=(x,y,z)\in\mathbb{R}^{N}$. $d$ is the head dimension. The upper triangle contains learnable $\textrm{LieRE}_{\theta}$ parameter, while the lower triangle contains their negatives, -$\textrm{LieRE}_{\theta}$, reflecting the skew-symmetric nature of the matrix.
[…]
@@
Figure 2: Comparisons of basis matrix $A$ of our work, Rope-Mixed [heo2024rotary], VisionLlama [chu2024visionllama]. First row: learnable parameter (yellow), not learnable parameter (blue) in basis matrix. Second row: rotation matrices $\exp(Ax)$ shared across the stem (gray) or different for each layer and head (colorful). MHA := Multi-Head-Attention, FF := Feed Forward, L[index] := Layer index, RxH := Rotation Matrix per Head, $\textrm{LieRE}_{\theta}$ := the number of LieRE parameters.
1.1 Contributions
We introduce Lie Relative Encodings (LieRE), a mechanism that allows the attention mechanism to learn how to utilize relative spatial information of its inputs. We show that LieRE is effective on both 2D and 3D inputs of various modalities. Beyond improving classification accuracy, LieRE also reduces the amount of compute and data required during training to achieve a fixed accuracy. On the CIFAR100 task, this translates to 3.5x fewer training steps to achieve the same accuracy as the DeiT baseline and outperforming the DeiT baseline trained on the full set with only a 70% subset of the data. Furthermore, LieRE is simple to implement and adaptable to modalities, requiring only a tokenizer that also outputs a position in $\mathbb{R}^{d}$ in addition to a standard embedding. In order to aid the reproducibility of our results we will post our code on github.
1.2 Related Work
1.2.1 Position Encodings
The fact that the original attention mechanism is invariant to the order of tokens has motivated the ongoing development of methods to incorporate positional information into the transformer architecture. We split our literature review into three broad classes of positional encodings: 1) absolute, 2) relative, and 3) contextual.
@@ 1.2.1 Position Encodings
Absolute encodings generally operate on a per token-level, modifying the embedding of a token to encode the location of the token in the intput. Methods such as sinusoidal and learned absolute encodings add vectors to the input token embedding [vaswani2017attention, devlin2019bert, dosovitskiy2020image]. Absolute position measures position with respect to an absolute reference, such as the start of the text or the top left corner of an image.
@@
Relative position encodings instead encode the relative positions of two tokens. One strategy is to learn an embedding for position deltas which can be incorporated into the attention mechanism [shaw2018self, liu2021swin, liu2022swinv2]. However, this incurs quadratic computational cost in terms of the number of tokens. Rotary Position Encodings (RoPE) avoid this cost by rotating the key and query vectors before the attention inner product. The algebraic properties of the block-diagonal rotation matrices used in RoPE ensures only relative positional information is captured in the attention mechanism [RoFormer2024Su]. RoPE is quite widely used in open source LLMs including the PaLM, Llama and Mixtral models [touvron2023llama, chowdhery2023palm, jiang2024mixtral]. However, RoPE can perform poorly on inference for larger context sizes than the model was trained on. This has spurred an active line of work extending RoPE to longer contexts, work which we review later.
@@
We refer to the last category of positional as contextual position embeddings. This category is defined by encodings that aim to capture semantic positional information lost in traditional absolute and relative position encodings, often motivated by reasoning or mathematical tasks. Contextual Position encodings achieve (CoPE) this by allowing the model to learn how the position is computed [golovneva2024contextual]. Abacus embeddings enable transformers to learn how to handle arithmetic by better exposing the digit structure of numbers [mcleish2024transformers].
1.2.2 Extensions of RoPE
@@
The efficiency and popularity of RoPE have led to several lines of work looking to extend it focused on specific domains.
@@
One notable one is context extension, which aims to address the fact that RoPE NLP models trained on short documents tend to perform poorly on long documents. Methods like NTK-aware context extension, YaRN and LongRoPE focus on enabling already trained models to handle long context, both with and without finetuning [ding2024longrope, peng2023yarn, tworkowski2024focused, chen2023extending].
@@
Another line of work has been specifically focused on adapting RoPE to image tasks. Both VisionLlama and RoPE-Mixed present relative position encodings inspired by RoPE that are able to encode 2D positional data [chu2024visionllama, heo2024rotary]. The primary difference is that RoPE-Mixed has a learnable component, whereas VisionLlama does not.
@@ 2.1 Lie Groups in the Context of Attention
In this section, we aim to provide a minimal introduction to Lie groups so that the reader is able to understand the mathematical motivations behind LieRE. Lie groups are well studied, especially in the context of representation theory, and standard texts including [fulton2013representation] are able to provide a more extensive introduction to the subject.
In this context, Lie groups are smooth sets of matrices that are closed under matrix multiplication and inversion. For every Lie group, the matrix exponential provides a smooth bijective map from a subset of $\mathbb{R}^{n\times n}$ (hereto referred to as the generator set) to the Lie group. The exponential map is a diffeomorphism and has the following key property for generators $U,V\in\mathbb{R}^{n\times n}$ close together:
@@
Section 3.1 will show that this is the key to being able to encode relative positions.
@@
Both RoPE (in the context of text) and RoPE-Mixed use block-diagonal rotation matrices with 2D rotations as blocks. These form a special Lie group that is commutative, allowing us to strengthen the statement in equation (1) to
$\exp(U-V)=\exp(U)\exp(V)^{-1}\boldsymbol{=}\exp(U)^{-1}\exp(V).$ (2)
@@
Our work examines the tradeoff between using the stronger property in equation (2) or increased capacity and the weaker property (1).
3 Methods
3.1 Rotations for Relative Positions in Attention
Before introducing the pseudocode we explain how LieRE modifies the attention mechanism and how this relates to Lie groups.
@@ 3.1 Rotations for Relative Positions in Attention
We start by using a learnable linear map to embed the positions as skew-symmetric matrices $P_{i}$ and $P_{j}$ and compute high dimensional rotations as $R_{i}=\exp(P_{i})$ and $R_{j}=\exp(P_{j})$. Learning in the space of skew symmetric matrices allows us to sidestep some of the difficulty that would come from learning on the manifold of rotation matrices.
Recall that, for every pair of tokens, the attention mechanism computes the inner products between their key and query vectors, $k_{i}^{T}q_{j}$, as shown in figure 1. We encode the positions by multiplying by the rotation matrices in the prior part. In particular, we update the keys and queries as $k_{i}^{\prime}=k_{i}R_{i}$ and $q_{j}=q_{j}R_{j}$. This results in an updated inner product of $(R_{i}k_{i})^{T}R_{j}q_{j}=k_{i}R_{i}^{T}R_{j}q_{j}=k_{i}R_{i}^{-1}R_{j}q_{j}.$
Recall that by equation 1, $R_{i}^{-1}R_{j}=\exp(P_{i})^{-1}\exp(P_{j})\approx\exp(P_{j}-P_{i})$. In other words, the inner product automatically computes the relative position encoding. Note that this is the mechanism behind the original RoPE paper. The difference is that LieRE uses the full set of high dimensional rotations and relaxes the constraint that equation 1 holds with equality. We make the description of the method more precise in the next section.
3.2 LieRE
LieRE modifies the attention mechanism by applying a rotation to the keys and queries before computing their inner products. However, where RoPE applies a block-diagonal rotation matrix with 2D rotation matrices as blocks, LieRE applies a $n\times n$ rotation matrix (Figure 1). Using the fact that every rotation matrix can be represented as the matrix exponential of a skew-symmetric matrix, we parametrize the rotations with generators before the matrix exponential. This skew-symmetric generator is computed via a learned linear map from the positions to the generator space. Let $x\in\mathbb{R}^{n}$ indicate the position of the token, and $A:\mathbb{R}^{d}\rightarrow\text{Skew}_{n}(\mathbb{R})$ be a learnable linear map to the set of skew-symmetric matrices. At the start of every forward pass, we compute the LieRE rotation matrix for each token as
$R_{LieRE}:=\exp(Ax).$
Then, for each attention mechanism, the keys and queries for a particular token are updated as $\mathrm{key^{\prime}}:=R_{LieRE}\cdot\mathrm{key},\mathrm{query^{\prime}}:=R_{LieRE}\cdot\mathrm{query}$ We present attention with LieRE and LieRE-Commute side by side in algorithms 1 and 2.
Algorithm 1 RoPE Attention (prior work for text), rot=rotated
1:procedure RoPE($X,p,d$)
2:     $\theta\leftarrow\frac{1}{10000^{2i/d}}\quad\forall i\in[0,d)$
3:     for $i\leftarrow 0$ to $d$ step $2$ do
4:         $X_{\text{rot}}^{i}\leftarrow X^{i}\cos p\theta^{i}-X^{i+1}\sin p\theta^{i}$
5:         $X_{\text{rot}}^{i+1}\leftarrow X^{i}\sin p\theta^{i}+X^{i+1}\cos p\theta^{i}$
6:     end for
7:     return $X_{\text{rotated}}$
8:end procedure
9:procedure RoPEAttention($Q,K,V$)
10:     $p\leftarrow\text{tokenPositions}$
[…]
12:     $K_{\text{rot}}\leftarrow\textsc{RoPE}(K,p,d)$
13:     $Q_{\text{rot}}\leftarrow\textsc{RoPE}(Q,p,d)$
14:     $\text{Attention}\leftarrow\text{softmax}\left(\frac{Q_{\text{rot}}K_{\text{rot}}^{T}}{\sqrt{d}}\right)V$
15:     return Attention
16:end procedure
Algorithm 2 LieRE Attention (our work), rot=rotated
1:procedure LieRE($X,p,A$)
[…]
4:     return $X_{\text{rot}}$
5:end procedure
6:procedure LieREAttention($Q,K,V,A$)
7:     $p\leftarrow\text{tokenPositions}$
8:     $K_{\text{rot}}\leftarrow\textsc{LieRE}(K,p,A)$
9:     $Q_{\text{rot}}\leftarrow\textsc{LieRE}(Q,p,A)$
10:     $\text{Attention}\leftarrow\text{softmax}\left(\frac{Q_{\text{rot}}K_{\text{rot}}^{T}}{\sqrt{\text{dim}(K)}}\right)V$
11:     return Attention
12:end procedure
Outside of the generator scaling experiments, we set the rotation dimension equal to the head dimension and apply rotation on a per-attention-head basis.
Note that if we apply sparsity structure that enforces 2x2 blocks we recover commutativity. We refer to this configuration and LieRE-Commute. If we then have separate learnable parameters for every layer and head, we recover RoPE-mixed [heo2024rotary].

### v3

# LieRE: Lie Rotational Positional Encodings

@@ Abstract
Transformer architectures rely on position encodings to capture token dependencies. Rotary Position Encoding (RoPE) has emerged as a popular choice in language models due to its efficient encoding of relative position information through key-query rotations. However, RoPE faces significant limitations beyond language processing: it is constrained to one-dimensional sequence data and, even with learnable phases, offers limited representational capacity. We address these challenges with Lie Relative Encodings (LieRE), which replaces RoPE’s block-2D rotation matrix with a learned, dense, high-dimensional rotation matrix of variable density. Through extensive evaluation on three image datasets across 2D and 3D classification tasks, LieRE achieves 2% relative improvement over state-of-the-art baselines on 2D tasks and 1.5% on 3D tasks, while demonstrating superior generalization to higher resolutions. Our implementation is computationally efficient, with results reproducible on 4 A100 GPUs in 30 minutes on CIFAR100, and we release our code to facilitate further research.
Machine Learning, ICML
@@ 1 Introduction
The attention mechanism, particularly within transformer architectures, has revolutionized machine learning across diverse domains. However, attention is inherently permutation—invariant—it cannot leverage the sequential order of its inputs as information. This fundamental limitation necessitates additional mechanisms to encode positional information, enabling models to capture sequential and spatial dependencies crucial for tasks ranging from natural language processing to image understanding (Vaswani et al., 2017).
This challenge has sparked extensive research into positional encodings, which inject order information into the otherwise order-agnostic attention mechanism. The field has evolved from early approaches using fixed sinusoidal positional embeddings (Vaswani et al., 2017) to more sophisticated learned embeddings (Shaw et al., 2018a; Devlin et al., 2019; Dosovitskiy et al., 2020). Recent advances have introduced increasingly dynamic methods, including relative position representations (Shaw et al., 2018b; Dosovitskiy et al., 2020) and rotary positional embeddings (Su et al., 2024; Heo et al., 2024), demonstrating the critical role of position encoding in enabling transformers to effectively process ordered sequences.
In particular, Rotary Position Encoding (RoPE) has emerged as an elegant solution for encoding relative positional information in transformers (Su et al., 2024). RoPE works by applying a rotation matrix to each token’s keys and queries, where the rotation angle depends on the token’s absolute position in the sequence. The key insight is that when two tokens interact through the attention inner product, their rotated representations naturally encode their relative distance. For example, when a token at position five attends to a token at position two, RoPE enables the model to understand that these tokens are three positions apart, regardless of their absolute positions in the sequence. This translation-invariant property makes RoPE particularly efficient at capturing position-dependent patterns in text, which has led to its adoption in popular open-source language models such as LLaMA and Mixtral.
@@
Figure 1: Overview of LieRE, where $A$ is a learnable skew symmetric matrix and $R_{i}=\exp(A[x\;y\;z])\in\mathbb{R}^{d\times d}$ is the rotation matrix for the first patch in the top left corner of the input with the position $P_{i}=(x,y,z)\in\mathbb{R}^{N}$. $d$ is the head dimension. The upper triangle contains learnable $\textrm{LieRE}_{\theta}$ parameter, while the lower triangle contains their negatives, -$\textrm{LieRE}_{\theta}$, reflecting the skew-symmetric nature of the matrix.
Despite RoPE’s success in sequential tasks (Touvron et al., 2023; Chowdhery et al., 2023), it faces several fundamental limitations. First, RoPE is designed for one-dimensional sequence data with an exponential decay in frequencies, making it suboptimal for spatial relationships in images or spatiotemporal patterns in videos where distant pixels may have strong correlations. Second, when handling higher dimensional data, the challenge compounds significantly–learned relative position encodings must capture exponentially many relative positions for $n$-dimensional data, as demonstrated in 2D image data where positions must be encoded both horizontally and vertically (Shaw et al., 2018a). Third, RoPE’s reliance on sparse, block-diagonal rotation matrices with handcrafted basis functions constrains its ability to learn complex spatial dependencies (Chu et al., 2024).
Our key insight is that Lie group theory provides a natural framework for generalizing relative position encodings to higher dimensions. We introduce Lie Relative Encodings (LieRE), which replaces RoPE’s rotation matrices with learned, dense rotations derived from Lie group generators. LieRE learns a basis of skew-symmetric matrices ${A_{i}}$, computes rotation matrices $R(p)=\exp(\sum_{i=0}^{n}p_{i}A_{i})$ for $n$-dimensional positions, and applies these rotations to keys and queries in the attention mechanism. The relative positions are then naturally captured through the inner product of the rotated keys and queries.
This approach addresses RoPE’s limitations in two key ways: (1) it handles higher-dimensional spatial relationships through Lie groups and (2) it increases representational capacity through variably dense, learned rotation matrices while requiring only 0.68% of parameters in a ViT-B model. Importantly, LieRE can be implemented with a single modification to existing architectures.
To assess the impact of LieRE and other positional encodings on ViT performance, we evaluate several encoding schemes across diverse tasks, including 2D and 3D image classification. Additionally, we investigate a fundamental spatial reasoning task where models must identify arrow directions. Despite their sophistication, contemporary multimodal LLMs struggle with this seemingly simple task. Our experiments reveal that successful completion of this task specifically requires relative position encodings, highlighting their crucial role in spatial understanding.
@@
Figure 2: Comparisons of basis matrix $A$ of our work, Rope-Mixed (Heo et al., 2024), VisionLlama (Chu et al., 2024). First row: learnable parameter (yellow), not learnable parameter (blue) in basis matrix. Second row: rotation matrices $\exp(Ax)$ shared across the stem (gray) or different for each layer and head (colorful). MHA := Multi-Head-Attention, FF := Feed Forward, L[index] := Layer index, RxH := Rotation Matrix per Head, $\textrm{LieRE}_{\theta}$ := the number of LieRE parameters.
2 Related Work
2.1 Position Encodings
We organize our literature review into three broad classes of positional encodings: 1) absolute, 2) relative, and 3) contextual.
@@ 2.1 Position Encodings
Absolute position in this context refers to a position with respect to a consistent reference, usually the start of a tex or top left corner of an image. Absolute encodings generally operate on a per token-level, modifying the embedding of a token to encode the location of the token in the input. Methods such as sinusoidal and learned absolute encodings directly add vectors to the input token embedding (Vaswani et al., 2017; Devlin et al., 2019; Dosovitskiy et al., 2020).
@@
Relative position encodings in contrast encode the relative positions of two tokens. One strategy is to learn an set of embeddings for position deltas which can be incorporated into the attention mechanism (Shaw et al., 2018b; Liu et al., 2021, 2022). However, this incurs quadratic computational cost in terms of the number of tokens as a separate embedding is required for every pair of tokens. Rotary Position Encodings (RoPE) avoid this cost by rotating the key and query vectors before the attention inner product. The algebraic properties of the block-diagonal rotation matrices used in RoPE ensures only relative positional information is captured in the attention mechanism (Su et al., 2024). RoPE is quite widely used in open source LLMs including the PaLM, Llama and Mixtral models (Touvron et al., 2023; Chowdhery et al., 2023; Jiang et al., 2024). However, RoPE can perform poorly on inference for larger context sizes than the model was trained on. This has spurred an active line of work extending RoPE to longer contexts, work which we review later.
@@
We refer to the last category of positional encodings as contextual position embeddings. This category is defined by encodings that aim to capture semantic positional information lost in traditional absolute and relative position encodings, often motivated by reasoning or mathematical tasks. Contextual Position encodings achieve (CoPE) this by allowing the model to learn how the position is computed (Golovneva et al., 2024). Abacus embeddings enable transformers to learn how to handle arithmetic by better exposing the digit structure of numbers (McLeish et al., 2024).
2.2 Extensions of RoPE
@@
The efficiency and popularity of RoPE have led to several lines of work building off of it.
@@
One notable one is context extension, which aims to address the fact that RoPE NLP models trained on short documents tend to perform poorly on long documents. Methods like NTK-aware context extension, YaRN and LongRoPE focus on enabling already trained models to handle long context, both with and without finetuning (Ding et al., 2024; Peng et al., 2023; Tworkowski et al., 2024; Chen et al., 2023).
@@
The final, and most relevant, line of work has been specifically focused on adapting RoPE to image tasks. Both VisionLlama and RoPE-Mixed present relative position encodings inspired by RoPE capable of encoding 2D positional data (Chu et al., 2024; Heo et al., 2024). The primary difference is that RoPE-Mixed has a learnable component, whereas VisionLlama does not.
@@ 3.1 Lie Groups in the Context of Attention
In this section, we aim to provide a minimal introduction to Lie groups so that the reader is able to understand the mathematical motivations behind LieRE. Lie groups are well studied, especially in the context of representation theory, and standard texts including (Fulton & Harris, 2013) are able to provide a more extensive introduction to the subject.
In this context, Lie groups are smooth sets of matrices that are closed under matrix multiplication and inversion. For every Lie group, the matrix exponential provides a smooth bijective map from a subset of $\mathbb{R}^{n\times n}$, also known as the Lie Algebra, to the Lie group. The exponential map is a diffeomorphism and has the following key property for $U,V\in\mathbb{R}^{n\times n}$ close together:
@@
Both RoPE (in the context of text) and RoPE-Mixed use block-diagonal rotation matrices with 2D rotations as blocks 4. These form a special Lie group that is commutative, allowing us to strengthen the statement in (1) to
$\exp(U-V)=\exp(U)\exp(V)^{-1}\boldsymbol{=}\exp(V)^{-1}\exp(U).$ (2)
@@
Our work examines the tradeoff between using the stronger property in (2) or increased capacity and the weaker property (1).
Figure 3: Pseudocode for the Lie Rotary Embedding (LieRE) attention mechanism. Unlike RoPE, LieRE allows for dense rotation matrices.
3.2 Attention Mechanism
LieRE is a modification of the standard attention mechanism to introduce positional information, which we review below. The modification we propose is independent of whether we use multiple heads, so we focus on single-headed attention for simplicity.
Let $X\in\mathbb{R}^{n\times d}$ be the set of input embeddings and $W_{Q},W_{K},W_{v}$ be learnable matrices. Let $Q=XW_{Q},K=XW_{K},V=XW_{V}$ be the keys, queries and values respectively. The outputs are computed as $\text{scores}=\frac{QK^{\top}}{\sqrt{d_{k}}}$, $\mathcal{W}=\text{softmax}(\text{scores})$ and final outputs $z=\mathcal{W}V$. We let $Q_{i}$ and $K_{i}$ denote the $i$th rows of $Q$ and $K$ respectively.
4 Method
LieRE is a simple modification to the attention mechanism that is presented in algorithm 3. Recall that we assume that positions are $n$-dimensional vectors, a matrix $A$ is skew-symmetric if $A^{T}=-A$, and that the matrix exponential of a skew-symmetric matrix, call it $A$, is always a high dimensional rotation matrix.
@@ 4 Method
When encoding positions $p\in\mathbb{R}^{n}$, LieRE learns a skew-symmetric basis of matrices $\{A_{i}\}$ for $i\in[n]$. It encodes a position by writing it in this basis, $\sum\limits_{i=0}^{n}p_{i}A_{i}$. We then map the resulting skew-symmetric matrix to a high-dimensional rotation via the matrix exponential. $R(p)=\exp\left(\sum\limits_{i=0}^{n}p_{i}A_{i}\right)$ (Figure 1). Learning in the space of skew-symmetric matrices allows us to sidestep some of the difficulty that would come from learning on the manifold of rotation matrices (Figure 2).
Figure 4: Pseudocode for the Rotary Position Embedding (RoPE) attention mechanism.
LieRE uses the rotation matrix computed above to modify the keys and queries of the standard attention mechanism. LieRE’s final step is to modify token $i$’s query and keys as $Q^{\prime}_{i}=R(p_{i})Q_{i}$ and $K^{\prime}_{i}=R(p_{i})K_{i}$. This modifies the score between tokens $i,j$ to be $X_{i}^{T}W_{Q}^{T}R(p_{i})^{T}R(p_{j})W_{K}X_{j}$. Recalling that $R^{T}=R^{-1}$ for any orthogonal matrix $R$ helps illustrate the encoding of relative positions in (1). Note that the only difference between LieRE and RoPE-Mixed is that the latter constrains the rotations to be block-diagonal with block size two.
We include the pseudocode for the LieRE attention in algorithm 3 in addition to the standard RoPE attention (algorithm 4). In practice, we compute the rotation matrices at the start of the forward pass. By default, the skew symmetric bases are learned separately for every layer and attention head except in the experimental section focused on sharing parameters across heads and layers.
Adjusting the skew-symmetric basis matrices’ block width allows us to incrementally adjust the capacity allocated towards position encoding. We specify the basis block width as a subscript, eg. LieRE8. When not specified, the block size is equal to the head dimension. If we set the block size to 2, we recover RoPE-Mixed (Heo et al., 2024). For a ViT-B model the head dimension is 64.